Caspian Sea Drinks is considering buying the J-Mix 2000. It will allow them to make and sell more product. The machine cost $1.56 million and create incremental cash flows of $649,261.00 each year for the next five years. The cost of capital is 9.75%. What is the profitability index for the J-Mix 2000?
Answer #1

Time Cash Flows Present Value Factor Present Value n A B = (1/1.0975)^n C = A*B 6,49,261.00 0.9112 5,91,581.7768 2 6,49,261.0